OpenStudy (anonymous):

HELP : Part (ii) (b) (i) Express (√6) cos θ + (√10) sin θ in the form R cos(θ − α ), where R > 0 and 0◦< α < 90◦. Give the value of α correct to 2 decimal places. (ii) Hence, in each of the following cases, find the smallest positive angle θ which satisfies the equation (a) (√6) cos θ + (√10) sin θ = −4, (b) (√6) cos1/2θ + (√10) sin1/2θ = 3

OpenStudy (anonymous):

@hartnn @thomaster

OpenStudy (atlas):

do you know the expansion of cos(theta-alpha) ?

OpenStudy (anonymous):

i just need help for ( ii) part (b)

OpenStudy (atlas):

is it theta/2 or 1/(2theta)

OpenStudy (anonymous):

\[4 \cos \left( \theta/2 - \alpha \right) = 3\] \[\alpha = 53.238..\]

OpenStudy (anonymous):

\[\left( \theta/2 - \alpha \right) = \cos^{-1} \frac{ 3 }{ 4 }\]

OpenStudy (atlas):

yeah you are right.............now substitute the values

OpenStudy (anonymous):

theta / 2 = cos inverse 3/4 + alpha... theta = 2 ( cos inverse 3/4 + alpha) = 189.2952

OpenStudy (anonymous):

but thats not the right answer :S

OpenStudy (atlas):

let me check a moment

OpenStudy (anonymous):

@satellite73

OpenStudy (atlas):

k i got it

OpenStudy (atlas):

theta - alpha/2 can be negative also

OpenStudy (anonymous):

its just : theta /2 not alpha...

OpenStudy (atlas):

sry.......theta/2 - alpha can be negative also....................you know cos ^-1(3/4) is also -41.4

OpenStudy (atlas):

try this value you will get your answer.........i expect the answer is 23.6

OpenStudy (anonymous):

yep but the answer should be : 21.7 ~_~

OpenStudy (anonymous):

(Y) ! :D

OpenStudy (anonymous):

but one more question : how do we know when to take negative for cos-1 ?

OpenStudy (anonymous):

why take negative of cos inverse rather than positive o.o ?

OpenStudy (atlas):

in this case you should check with both to see which one is giving you the smallest angle.........because the question says find smallest positive angle

OpenStudy (anonymous):

suppose the question was to find all values of theta then i should take cos both positive and neg ?

OpenStudy (atlas):

yup.......and if the question is to find ALL the values of theta, then there are several answers.....remember trigonometric functions start repeating again and again

OpenStudy (anonymous):

:D ty :D !